# A modern perspective on rational homotopy theory

### v1

@@ 4.1. Fibrant replacement of $1$-reduced simplicial sets
We invoke Corollary 3.19 to produce functorial fibrant replacements in ${\mathcal{SS}\mathrm{ets}}_{1}$. Let $X$ be a $1$-reduced simplicial set. The geometric realization-singular complex adjunction from Proposition 2.7 provides the natural map
$X\rightarrow\operatorname{Sing}(|X|),$
which maps into the $1$st Eilenberg subcomplex $E_{1}\operatorname{Sing}(|X|)$. Thus, we obtain the commutative diagram
${X}$${{E_{1}\operatorname{Sing}(|X|)}}$${{\operatorname{Sing}(|X|),}}$
@@
is a weak homotopy equivalence, thus a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$. To do so, it suffices to show that the other two maps in the commutative diagram are weak homotopy equivalences. Firstly, Part (1) of the following theorem informs us that the natural map
$X\rightarrow\operatorname{Sing}(|X|)$
is a weak homotopy equivalence; Part (2) is for later use in this argument.
Theorem 4.1 ([GJ99, Chapter I, Theorem 11.4]).
@@
    For every simplicial set $X$, the natural map
    $X\rightarrow\operatorname{Sing}(|X|)$
    is a weak homotopy equivalence.
@@
    For every topological space $Y$, the natural map
    $|\operatorname{Sing}(Y)|\rightarrow Y$
    is a weak homotopy equivalence.
@@
Part (1) informs us that, for a simplicial set $X$, the natural map
$X\rightarrow\operatorname{Sing}(|X|)$
is a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}$, as $\operatorname{Sing}(|X|)$ is a Kan complex by Example 3.2.
As for the $1$st Eilenberg subcomplex inclusion
$E_{1}\operatorname{Sing}(|X|)\rightarrow\operatorname{Sing}(|X|),$
it is also a weak homotopy equivalence by a special case of the following theorem.
Theorem 4.3 ([May67, Theorem 8.4(iii)]).
If $X$ is a pointed Kan complex, then the $1$st Eilenberg subcomplex inclusion $E_{1}(X)\rightarrow X$ induces, for every $n\geq 2$, the isomorphism
$\pi_{n}\left(E_{1}(X)\right)\xrightarrow{\cong}\pi_{n}(X).$
In particular, if $X$ is a pointed and $1$-connected Kan complex, then the inclusion
$E_{1}(X)\rightarrow X$
is a weak homotopy equivalence.
Remark 4.4.
In other words, if $X$ is a pointed Kan complex, then the $1$st Eilenberg subcomplex inclusion
$E_{1}(X)\rightarrow X$
is a $1$-connected cover of $X$.
As an aside, using the construction of the homotopy groups of Kan complexes in [GJ99, Chapter I.7] together with Lemma 2.22, we can provide an alternative proof of Theorem 4.3 from that in [May67, Theorem 8.4(iii)].
@@
Let $X$ be a pointed Kan complex. Let $n\geq 2$. We first show that $\pi_{n}\left(E_{1}(X)\right)\rightarrow\pi_{n}(X)$ is an epimorphism. As $X$ is a Kan complex, every element of $\pi_{n}(X)$ is of the form $[\omega]$ for a map
$\omega\colon S_{\mathrm{simp}}^{n}\rightarrow X.$
As $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is $1$-reduced, so the universal property of the $1$st Eilenberg subcomplex from Lemma 2.22 yields the commutative diagram
@@
It remains to show that $\pi_{n}\left(E_{1}(X)\right)\rightarrow\pi_{n}(X)$ is a monomorphism. Given a pointed homotopy
$H\colon S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}\rightarrow X,$
we observe that $S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}$ is $1$-reduced by Lemma 2.15 because $S_{\mathrm{simp}}^{n}$ is $1$-reduced, so the universal property of the $1$st Eilenberg subcomplex from Lemma 2.22 yields the commutative diagram
@@
If $Y$ is a pointed and $1$-connected topological space, then the inclusion
$E_{1}\operatorname{Sing}(Y)\rightarrow\operatorname{Sing}(Y)$
is a weak homotopy equivalence.
@@
By Example 3.2, $\operatorname{Sing}(Y)$ is a Kan complex. By Theorem 4.1(2), $\operatorname{Sing}(Y)$ is $1$-connected. ∎
@@
Let $X$ be a $1$-reduced simplicial set. Then, the natural map
$X\rightarrow E_{1}\operatorname{Sing}(|X|)$
is a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$.
4.2. Generating sets of cofibrations of $1$-reduced simplicial sets
Our left transfer from Theorem 3.17 provides us with a combinatorial model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$, but not with explicit generating sets of cofibrations and acyclic cofibrations. As a first remedy, we produce an explicit generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@ 4.2. Generating sets of cofibrations of $1$-reduced simplicial sets
Let $I=\left\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\right\}$ denote the generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}$. Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. The following statements are equivalent.
@@
    ${R}_{1}(I)=\left\{{R}_{1}(i_{n})\colon{R}_{1}(\partial\Delta[n])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 0\right\}.$
@@
Firstly, we show that (1) implies (2). If $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}$, then $f$ has the right lifting property with respect to all monomorphisms in ${\mathcal{SS}\mathrm{ets}}$. In particular, $f$ has the right lifting property with respect to all monomorphisms in ${\mathcal{SS}\mathrm{ets}}_{1}$, so $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Then, (2) implies (3) as, by Lemma 3.14, ${R}_{1}(I)$ is a set of cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Lastly, we show that (3) implies (1). Suppose $f$ has the right lifting property with respect to ${R}_{1}(I)$. To prove that $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}$, we solve, for $n\geq 0$, a lifting problem
${{\partial\Delta[n]}}$${X}$${{\Delta[n]}}$${{Y.}}$$\scriptstyle{i_{n}}$$\scriptstyle{f}$
Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, we equivalently solve a lifting problem
${{{R}_{1}(\partial\Delta[n])}}$${X}$${{{R}_{1}(\Delta[n])}}$${{Y.}}$$\scriptstyle{{R}_{1}(i_{n})}$$\scriptstyle{f}$
@@
Even though our treatment of Proposition 4.7 is biased towards the equivalence of (2) and (3), which grants us our generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$, we would be remiss if we neglected the equivalence of (1) and (2). The fact that (1) and (2) are equivalent means that ${\mathcal{SS}\mathrm{ets}}_{1}$ has no pathological acyclic fibrations. As we shall see later in this section, Proposition 4.7 finds no analog in the study of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$. The reason is that ${\mathcal{SS}\mathrm{ets}}_{1}$ has fibrations that are dissimilar to Kan fibrations, a phenomenon that we also investigate in this section. Thus, the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ is more than a restriction of the model category ${\mathcal{SS}\mathrm{ets}}$ to $1$-reduced simplicial sets; after all, the model category axioms are generally not amenable to such a restriction. Instead, the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ formalizes a different homotopy theory that is interesting in its own right.
@@ 4.3. Acyclic cofibrations of $1$-reduced simplicial sets
Our attempt to emulate Proposition 4.7 towards a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ encounters a problem: the $1$-reduction of a boundary inclusion is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, but the $1$-reduction of a horn inclusion need not be an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@
For $n\geq 1$ and $0\leq k\leq n$, let $j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]$ denote the horn inclusion in ${\mathcal{SS}\mathrm{ets}}$. Then, ${R}_{1}\left(j_{n,k}\right)$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $n\neq 2$.
@@
Nevertheless, we can partially work our way around this obstacle. Let
$J=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and}0\leq k\leq n\right\}$
denote the generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}$. Write
${R}_{1}(J)=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 1\textrm{ and }0\leq k\leq n\right\},$
which is a set of cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ by Lemma 3.14, but not a set of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ by Lemma 4.8. To extract a candidate generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}$, we use Lemma 3.13 to strike out
@@
    the maps ${R}_{1}(j_{2,0})$, ${R}_{1}(j_{2,1})$, and ${R}_{1}(j_{2,2})$ which are not weak homotopy equivalences.
Therefore, we write
$\widetilde{J}=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 3\textrm{ and }0\leq k\leq n\right\},$
whose elements are acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}$, in order to write
${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\},$
which is a set of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$. Hence, having the right lifting property with respect to the set ${R}_{1}(\widetilde{J})$ is necessary for a map of $1$-reduced simplicial sets to be a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. We partially address the question of sufficiency. Namely, we establish that the set ${R}_{1}(\widetilde{J})$ detects fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@
    The unique map $c\colon X\rightarrow\Delta[0]$ has the right lifting property with respect to the set
    ${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\}.$
Hence, the set ${R}_{1}(\widetilde{J})$ detects fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@
Firstly, (1) and (2) are equivalent by the characterization of fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$ in Theorem 3.17.
If (2) holds, then $c\colon X\rightarrow\Delta[0]$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. Thus, $c\colon X\rightarrow\Delta[0]$ has the right lifting property with respect to the set ${R}_{1}(\widetilde{J})$, since every map in ${R}_{1}(\widetilde{J})$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. Therefore, (3) holds.
Lastly, suppose (3) holds. We prove that (1) holds. For $n\geq 1$ and $0\leq k\leq n$, a lifting problem
${{\Lambda[n,k]}}$${X}$${{\Delta[n]}}$$\scriptstyle{h}$$\scriptstyle{j_{n,k}}$$\scriptstyle{g}$
is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, to a lifting problem
${{{R}_{1}\left(\Lambda[n,k]\right)}}$${X}$${{{R}_{1}\left(\Delta[n]\right).}}$$\scriptstyle{\eta}$$\scriptstyle{{R}_{1}\left(j_{n,k}\right)}$$\scriptstyle{\psi}$
In light of Lemma 3.13, we split our solution in three separate cases. If $n=1$, then Lemma 3.13(1) says that ${R}_{1}\left(j_{1,k}\right)$ is the identity map of the point, so our lifting problem can be solved by setting $\psi=\eta$. If $n\geq 3$, then our lifting problem can be solved, as we are assuming that $c\colon X\rightarrow\Delta[0]$ has the right lifting property with respect to the set ${R}_{1}(\widetilde{J})$. Lastly, if $n=2$, then Lemma 3.13(2) says that our lifting problem is
@@
where $x_{0}$ denotes the unique 00-simplex of $X$, so we take $\psi$ to be the constant map at $x_{0}$. ∎
@@
Proposition 4.10 ([Qui69, Part II, Proposition 2.12]).
Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. Suppose that $Y$ is a Kan complex. The following statements are equivalent.
  1. (1)
    The map $f\colon X\rightarrow Y$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
  2. (2)
    ${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\}.$
Hence, the set ${R}_{1}(\widetilde{J})$ detects fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ whose target is a Kan complex.
The additional fibrancy condition on the target is crucially used in Quillen’s proof, which suggests that ${R}_{1}(\widetilde{J})$ is not a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$. In his discussion of [Qui69, Part II, Proposition 2.12], Quillen affirms this suggestion and provides a hint towards showing that the fibrancy condition on the target in Proposition 4.10 is necessary. However, to the best of our knowledge, the question of whether ${R}_{1}(\widetilde{J})$ is a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ remains open. Our various attempts at producing a map in ${\mathcal{SS}\mathrm{ets}}_{1}$ that has the right lifting property with respect to ${R}_{1}(\widetilde{J})$ but is not a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ have been unsuccessful in the same fashion: our candidate maps are not fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$, but they also fail to have the right lifting property with respect to ${R}_{1}(\widetilde{J})$. We pose the following conjecture to the community.
@@
${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\}$
@@
4.4. Fibrations of $1$-reduced simplicial sets
The issues that we ran into in our study of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ suggest that there exist fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ that are dissimilar to Kan fibrations. To investigate this phenomenon, we characterize when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration. In fact, we produce two results to that end, which help us find an explicit example of a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ that is not a Kan fibration.
To begin with, we note that every Kan fibration of $1$-reduced simplicial sets is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.

### v2

@@ 4. Fibrations of $1$-reduced simplicial sets
The existence of such counterintuitive fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ serves not only as an impediment toward finding a generating set of acyclic cofibrations, but also as evidence that the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ should not be thought of as a restriction of the model structure on pointed simplicial sets. Instead, ${\mathcal{SS}\mathrm{ets}}_{1}$ encodes its own homotopy theory that has its individual intricacies, some of which we shed light on in this section.
@@ 4.1. Fibrant replacement of $1$-reduced simplicial sets
We invoke Corollary 3.18 to produce functorial fibrant replacements in ${\mathcal{SS}\mathrm{ets}}_{1}$. Let $X$ be a $1$-reduced simplicial set. The geometric realization-singular complex adjunction provides the map $X\rightarrow\operatorname{Sing}(|X|)$, which lands into $E_{1}\operatorname{Sing}(|X|)$. We get the commutative diagram
[…]
@@
is a weak equivalence, thus a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$. To do so, it suffices to show that the other two maps in the commutative diagram are weak equivalences. Firstly, part (1) of the following theorem says that $X\rightarrow\operatorname{Sing}(|X|)$ is a weak equivalence; part (2) is for later use in our argument.
Theorem 4.1 ([23, Chapter I, Theorem 11.4]).
@@
    For every simplicial set $X$, the map $X\rightarrow\operatorname{Sing}(|X|)$ is a weak equivalence.
@@
    For every topological space $Y$, the map $|\operatorname{Sing}(Y)|\rightarrow Y$ is a weak homotopy equivalence.
@@
By Example 3.2, part (1) defines a functorial fibrant replacement in $\mathcal{SS}\mathrm{ets}$.
The inclusion $E_{1}\operatorname{Sing}(|X|)\rightarrow\operatorname{Sing}(|X|)$ is also a weak equivalence by the following theorem.
Theorem 4.3 ([30, Theorem 8.4(iii)]).
If $X$ is a pointed Kan complex, then the 11st Eilenberg subcomplex inclusion $E_{1}(X)\rightarrow X$ induces, for every $n\geq 2$, an isomorphism $\pi_{n}(E_{1}(X))\xrightarrow{\cong}\pi_{n}(X)$. In particular, if $X$ is a pointed and $1$-connected Kan complex, then the inclusion $E_{1}(X)\rightarrow X$ is a weak equivalence.
In other words, if $X$ is a pointed Kan complex, then the inclusion $E_{1}(X)\rightarrow X$ is a $1$-connected cover of $X$. As an aside, using the construction of the homotopy groups of Kan complexes from [23, Chapter I.7] together with Proposition 2.18, we give an alternative proof of Theorem 4.3 to that in [30, Theorem 8.4(iii)].
@@
We first show that $\pi_{n}(E_{1}(X))\rightarrow\pi_{n}(X)$ is an epimorphism. As $X$ is a Kan complex, every element of $\pi_{n}(X)$ is of the form $\omega]$ for a map $\omega\colon S_{\mathrm{simp}}^{n}\rightarrow X$, and the universal property of the 11st Eilenberg subcomplex yields the commutative diagram
@@
It remains to show that $\pi_{n}(E_{1}(X))\rightarrow\pi_{n}(X)$ is a monomorphism. Given a homotopy $H\colon S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}\rightarrow X$, we observe that $S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}$ is $1$-reduced by Lemma 2.11, so we obtain the commutative diagram
@@
We infer that the inclusion $E_{1}\operatorname{Sing}(|X|)\rightarrow\operatorname{Sing}(|X|)$ is also a weak equivalence.
@@
If $Y$ is a pointed and $1$-connected space, then $E_{1}\operatorname{Sing}(Y)\rightarrow\operatorname{Sing}(Y)$ is a weak equivalence.
@@
By Example 3.2, $\operatorname{Sing}(Y)$ is a Kan complex. By Theorem 4.1(2), $\operatorname{Sing}(Y)$ is 11-connected. ∎
@@
Let $X$ be a $1$-reduced simplicial set. Then, the natural map $X\rightarrow E_{1}\operatorname{Sing}(|X|)$ is a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$.
4.2. Generating cofibrations of $1$-reduced simplicial sets
Our left transfer from Theorem 3.17 provides us with a combinatorial model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$, but not with explicit generating sets of cofibrations and acyclic cofibrations. As a first remedy, we extract an explicit generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ from the set of boundary inclusions for $\mathcal{SS}\mathrm{ets}$} using the $1$-reduction functor $R_{1}$.
@@ 4.2. Generating cofibrations of $1$-reduced simplicial sets
Let $I=\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\}$ denote the generating set of cofibrations for $\mathcal{SS}\mathrm{ets}$}. Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. The following statements are equivalent.
@@
    ${R}_{1}(I)=\{{R}_{1}(i_{n})\colon{R}_{1}(\partial\Delta[n])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 0\}.$
@@
Firstly, we show that (1) implies (2). If $f$ is an acyclic fibration in $\mathcal{SS}\mathrm{ets}$}, then $f$ has the right lifting property with respect to all monomorphisms in $\mathcal{SS}\mathrm{ets}$}. In particular, $f$ has the right lifting property with respect to all monomorphisms in ${\mathcal{SS}\mathrm{ets}}_{1}$, so $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. Also, (2) implies (3) as, by Lemma 3.14, ${R}_{1}(I)$ is a set of cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$. Lastly, we show that (3) implies (1). Suppose $f$ has the right lifting property with respect to ${R}_{1}(I)$. To prove that $f$ is an acyclic fibration in $\mathcal{SS}\mathrm{ets}$}, we solve, for $n\geq 0$, a lifting problem
[…]
Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, we equivalently solve a lifting problem
[…]
@@
Even though our treatment of Proposition 4.6 is biased toward the equivalence of statements (2) and (3), which grants us our generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$, we would be remiss if we neglected the equivalence of statements (1) and (2). The fact that statements (1) and (2) are equivalent means that ${\mathcal{SS}\mathrm{ets}}_{1}$ has no pathological acyclic fibrations, which follows from the fact that ${\mathcal{SS}\mathrm{ets}}_{1}$ has no pathological cofibrations. As we shall see later in this section, Proposition 4.6 finds no analog in the study of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ because ${\mathcal{SS}\mathrm{ets}}_{1}$ has fibrations that are dissimilar to Kan fibrations.
@@ 4.3. Acyclic cofibrations of $1$-reduced simplicial sets
Our attempt to emulate Proposition 4.6 toward a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ encounters a problem: the $1$-reduction of a boundary inclusion is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, but the $1$-reduction of a horn inclusion need not be an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@
For $n\geq 1$ and $0\leq k\leq n$, let $j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]$ denote the horn inclusion in $\mathcal{SS}\mathrm{ets}$}. Then, ${R}_{1}(j_{n,k})$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $n\neq 2$.
@@
We can partially work our way around this obstacle. Let $J=\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and }0\leq k\leq n\}$ denote the generating set of acyclic cofibrations for $\mathcal{SS}\mathrm{ets}$}. Write
${R}_{1}(J)=\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 1\textrm{ and }0\leq k\leq n\}.$
To extract a candidate generating set of acyclic cofibrations for $\mathcal{SS}\mathrm{ets}$}, we use Lemma 3.13 to strike out
@@
    the maps ${R}_{1}(j_{2,0})$, ${R}_{1}(j_{2,1})$, and ${R}_{1}(j_{2,2})$ which are not weak equivalences.
Therefore, we write $\widetilde{J}=\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 3\textrm{ and }0\leq k\leq n\}$, whose elements are acyclic cofibrations in $\mathcal{SS}\mathrm{ets}$}, in order to write
${R}_{1}(\widetilde{J})=\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\},$
[…]
@@
    The map $X\rightarrow\Delta[0]$ has the right lifting property with respect to the set
    ${R}_{1}(\widetilde{J})=\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\}.$
[…]
@@
[…]
is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, to a lifting problem
[…]
where $x_{0}$ denotes the unique 0-simplex of $X$, so we take $\psi$ to be the constant map at $x_{0}$. ∎
@@
Proposition 4.9 ([33, Part II, Proposition 2.12]).
[…]
@@
${R}_{1}(\widetilde{J})=\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\}$
@@
4.4. Pathological fibrations of $1$-reduced simplicial sets
The issues that we ran into in our study of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ suggest that there exist fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ that are dissimilar to Kan fibrations. To investigate this phenomenon, we characterize when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration. In fact, we produce two results that help us find an explicit example of a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ that is not a Kan fibration. To begin with, we note that every Kan fibration of $1$-reduced simplicial sets is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.